# Unforgettable Lessons from Forgettable Images: Intra-Class Memorability Matters in Computer Vision Tasks

### v1

@@ 5 Conclusion and Future Works
Previous studies typically present participants with sequences of images where each sequence contains images from multiple categories. However, image category can significantly influence memorability [7, 15, 30, 34, 5]. For example, dog images are generally more memorable than mountain images [30]. To eliminate the influence of image categories on memorability, we design a new human behavior experiment focusing on intra-class memorability. Fig. S1 compares our experimental protocol with previous approaches [7, 26, 13, 21, 15]. Unlike previous methods that mix multiple categories in a single sequence, we separate image sequences by category. Each participant views image sequences from a single category, with breaks in between image sequences from different categories to reduce category interference on memorability.
@@
Figure S1: The differences between our experimental protocol and previous studies. Our experimental design differs from previous studies [7, 26, 13, 21, 15]. In Row 1, “Others” represents the protocol used in previous studies, where participants are exposed to images from multiple categories within a single sequence. In Row 2, “Ours” represents our protocol, which requires participants to view only images from a single category per sequence, with the dotted line indicating the break between sequences. Different colored squares represent images from distinct categories, and the numbers indicate the position of each image within its category (e.g., B2 is the second image in category B). The black ”X” represents a foil image that appears only once. In our protocol, the foil image ”X” belongs to the same category as the target images in the same sequence.
S2 The Effect of Time Interval $t$ in Human Behavior Experiments
@@
In Sec. 3.2, we state that human memory retention tends to weaken over longer intervals. That is, a longer interval $t$ makes correct identification of the repeated images more challenging. See Sec. 3.1 for the definition of interval $t$. We demonstrate this effect in Fig. S2. Fig. S2 shows that identification accuracy of the repeated images and F1 score decrease with the longer interval $t$. Additionally, we conducted an experiment with $t$ exceeding 32, during which most participants struggled to complete the task. Therefore, we empirically set $t\in\{8,16,24,32\}$ in the actual experiments.
@@
Figure S2: Human behavior experiment scores. This figure presents human performance at different intervals of $t$ (see Sec. 3.2 for the definition of $t$). The x-axis represents $t$, and the y-axis represents the score, with the red line indicating the accuracy score and the green line indicating the F1 score [53]. The values represent the average accuracy and F1 scores [53], computed across all trials with the same $t$. The error bars represent the standard error.
S3 Correlation Analysis Between Feature Distance and ICMscores
@@
To investigate the relationship between an image’s feature distance to its cluster center and its ICMscore, we use a ResNet [20] model pre-trained on ImageNet1k [11] to extract feature representations. For each category, the cluster center is defined as the mean of all image features in that category. We compute the Pearson correlation between each feature’s L2 distance to the cluster center and the ICMscore.
@@
The results, presented in Fig. S3, reveal a positive correlation between feature distance and ICMscore across three object classes: teapot, dog, and jellyfish. Images with features further away from their category cluster center tend to be more memorable.
@@
Figure S3: Correlation analysis between the feature distances to cluster center from individual images and their corresponding ICMscore. Each image in the ICMD dataset is represented as a scatter point, with its corresponding ICMscore shown on the y-axis and the L2 distance from their image features to their categorical centriods on the x-axis. The panels from the left to the right represent the scatter plots based on the three classes, “teapot”, “dog”, and “jellyfish”. The dashed lines, derived from linear regression, indicate a positive correlation between feature distance and ICMscore. The Pearson correlation values $r$ for “teapot,” “dog,” and “jellyfish” are indicated in the subplot titles: 0.26, 0.25, and 0.21.
S4 Confusion Matrices in Image Recognition Experiments
@@
Following the image recognition experiment described in Sec. 4.2, we observe that models trained on LM images perform better in image recognition tasks than HM and MM images. To further demonstrate this difference, we compare the confusion matrices of models trained on three subsets. As shown in Fig. S4, the model trained on the LM images has a lower error rate across most classes, with the largest improvement in the “teapot” class, where its accuracy increases by 17% and 16% in comparison to the HM and MM models.
@@
Figure S4: Accuracy confusion matrices for models trained on HM, MM, and LM images classified based on human performances in behavioral experiments. Each confusion matrix shows the proportion of classifying images as predicted labels (columns) given the ground truth (rows). Values in all the entries of the matrices represent the average classification proportion over 5 runs. See the color bar on the right for the classification proportion.
@@
S5 Performance of Models Using Memorability Predictors from the Previous Works
@@
We replace ICMscore in ICMD with the memorability score from prior work [26] and split the dataset into two subsets (highly memorable images and low memorable images) following the same way in Sec. 4.2. In comparison, we also use our ICMscore predictor, proposed in Sec. 4.1, to divide the ICMD dataset into three subsets LM, MM, and HM. The three subsets are then used to train image recognition models and continual learning models during replays.
@@
Fig. S6 shows image recognition performance across the LM-model, the MM-model, and the HM-model. See Sec. 4.2 for the result discussion. In Fig. S6, we present the continual learning performance of various models with LM, MM, and HM images for replays. See Sec. 4.3 for the result discussion.
@@
Figure S5: Performance comparison in Top-1 Accuracy between image classification models trained on HM, MM, and LM images using different memorability predictors. (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in [26]. Within each subplot, from the top to the bottom, top-1 accuracy is reported for models trained on LM images (blue), HM images (green), and MM images (yellow). The chance is 20% for the 5-way classification (dash lines produced by the random model in red). The error bar is the standard error calculated over 5 independent runs with different random seeds.
@@
Figure S6: Continual learning performance of AI models using HM, MM, and LM images for replays (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in [26]. Within each subplot, from left to right, the performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays is presented in Acc (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for the introduction to the evaluation metrics and the continual learning models. The error bars are the standard errors calculated over 5 independent runs with different random seeds.

### v2

@@ A.1 Difference in Human Behavior Experiments from the Existing Literature
Previous studies typically present participants with sequences of images where each sequence contains images from multiple categories. However, image category can significantly influence memorability (Bylinskii et al., 2015; Goetschalckx & Wagemans, 2019; Kramer et al., 2023; Lu et al., 2020; Borkin et al., 2013). For example, dog images are generally more memorable than mountain images (Kramer et al., 2023). To eliminate the influence of image categories on memorability, we design a new human behavior experiment focusing on intra-class memorability. Fig. A1 compares our experimental protocol with previous approaches (Bylinskii et al., 2015; Khosla et al., 2015; Dubey et al., 2015; Hovhannisyan et al., 2021; Goetschalckx & Wagemans, 2019). Unlike previous methods that mix multiple categories in a single sequence, we separate image sequences by category. Each participant views image sequences from a single category, with breaks in between image sequences from different categories to reduce category interference on memorability.
@@
Figure A1: The differences between our experimental protocol and previous studies. Our experimental design differs from previous studies (Bylinskii et al., 2015; Khosla et al., 2015; Dubey et al., 2015; Hovhannisyan et al., 2021; Goetschalckx & Wagemans, 2019). In Row 1, “Others” represents the protocol used in previous studies, where participants are exposed to images from multiple categories within a single sequence. In Row 2, “Ours” represents our protocol, which requires participants to view only images from a single category per sequence, with the dotted line indicating the break between sequences. Different colored squares represent images from distinct categories, and the numbers indicate the position of each image within its category (e.g., B2 is the second image in category B). The black ”X” represents a foil image that appears only once. In our protocol, the foil image ”X” belongs to the same category as the target images in the same sequence.
A.2 The Effect of Time Interval $t$ in Human Behavior Experiments
@@
In Sec. 3.2, we state that human memory retention tends to weaken over longer intervals. That is, a longer interval $t$ makes correct identification of the repeated images more challenging. See Sec. 3.1 for the definition of interval $t$. We demonstrate this effect in Fig. A2. Fig. A2 shows that identification accuracy of the repeated images and F1 score decrease with the longer interval $t$. Additionally, we conducted an experiment with $t$ exceeding 32, during which most participants struggled to complete the task. Therefore, we empirically set $t\in\{8,16,24,32\}$ in the actual experiments.
@@
Figure A2: Human behavior experiment scores. This figure presents human performance at different intervals of $t$ (see Sec. 3.2 for the definition of $t$). The x-axis represents $t$, and the y-axis represents the score, with the red line indicating the accuracy score and the green line indicating the F1 score (Yacouby & Axman, 2020). The values represent the average accuracy and F1 scores (Yacouby & Axman, 2020), computed across all trials with the same $t$. The error bars represent the standard error.
A.3 Correlation Analysis Between Feature Distance and ICMscores
@@
To investigate the relationship between an image’s feature distance to its cluster center and its ICMscore, we use a ResNet (He et al., 2015) model pre-trained on ImageNet1k (Deng et al., 2009) to extract feature representations. For each category, the cluster center is defined as the mean of all image features in that category. We compute the Pearson correlation between each feature’s L2 distance to the cluster center and the ICMscore.
@@
The results, presented in Fig. A3, reveal a positive correlation between feature distance and ICMscore across three object classes: teapot, dog, and jellyfish. Images with features further away from their category cluster center tend to be more memorable.
@@
Figure A3: Correlation analysis between the feature distances to cluster center from individual images and their corresponding ICMscore. Each image in the ICMD dataset is represented as a scatter point, with its corresponding ICMscore shown on the y-axis and the L2 distance from their image features to their categorical centriods on the x-axis. The panels from the left to the right represent the scatter plots based on the three classes, “teapot”, “dog”, and “jellyfish”. The dashed lines, derived from linear regression, indicate a positive correlation between feature distance and ICMscore. The Pearson correlation values $r$ for “teapot,” “dog,” and “jellyfish” are indicated in the subplot titles: 0.26, 0.25, and 0.21($p<0.05$).
A.4 Confusion Matrices in Image Recognition Experiments
@@
Following the image recognition experiment described in Sec. 4.2, we observe that models trained on LM images perform better in image recognition tasks than HM and MM images. To further demonstrate this difference, we compare the confusion matrices of models trained on three subsets. As shown in Fig. A4, the model trained on the LM images has a lower error rate across most classes, with the largest improvement in the “teapot” class, where its accuracy increases by 17% and 16% in comparison to the HM and MM models.
@@
Figure A4: Accuracy confusion matrices for models trained on HM, MM, and LM images classified based on human performances in behavioral experiments. Each confusion matrix shows the proportion of classifying images as predicted labels (columns) given the ground truth (rows). Values in all the entries of the matrices represent the average classification proportion over 5 runs. See the color bar on the right for the classification proportion.
@@
A.5 Performance of Models Using Memorability Predictors from the Previous Works
@@
We replace ICMscore in ICMD with the memorability score from prior work (Khosla et al., 2015) and split the dataset into two subsets (highly memorable images and low memorable images) following the same way in Sec. 4.2. In comparison, we also use our ICMscore predictor, proposed in Sec. 4.1, to divide the ICMD dataset into three subsets LM, MM, and HM. The three subsets are then used to train image recognition models and continual learning models during replays.
@@
Fig. A6 shows image recognition performance across the LM-model, the MM-model, and the HM-model. See Sec. 4.2 for the result discussion. In Fig. A6, we present the continual learning performance of various models with LM, MM, and HM images for replays. See Sec. 4.3 for the result discussion.
@@
Figure A5: Performance comparison in Top-1 Accuracy between image classification models trained on HM, MM, and LM images using different memorability predictors. (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in (Khosla et al., 2015). Within each subplot, from the top to the bottom, top-1 accuracy is reported for models trained on LM images (blue), HM images (green), and MM images (yellow). The chance is 20% for the 5-way classification (dash lines produced by the random model in red). The error bar is the standard error calculated over 5 independent runs with different random seeds.
@@
Figure A6: Continual learning performance of AI models using HM, MM, and LM images for replays (a) shows results using our ICMscore predictor from Sec. 4.1, while (b) presents results from the regressor proposed in (Khosla et al., 2015). Within each subplot, from left to right, the performance of continual learning models using HM (green), MM (yellow), and LM (blue) images for replays is presented in Acc (left), BWT (middle), and Forgetting (right). See Sec. 4.3 for the introduction to the evaluation metrics and the continual learning models. The error bars are the standard errors calculated over 5 independent runs with different random seeds.
A.6 Performance of Models Using ICMscore without Time Penalty
We modify ICMscore by removing the time penalty, defining $s_{i}^{\prime}=1$ (correct) and $s_{i}^{\prime}=0$ (incorrect), while keeping ICMscore(I)$=\frac{\sum_{i=1}^{n}s_{i}}{n}$. Following the dataset splitting protocol from Sec. 4.2, we train image recognition and continual learning models on LM, MM, and HM subsets.
Fig. 7(a) shows that removing the time penalty reduces recognition best performance by 2% compared to Sec. 4.2. Similarly, the best continual learning performance in Fig. 7(b) degrades by 3% versus Sec. 4.3, narrowing the gap between memorability groups. Notably, the key finding remains consistent: the LM subset helps train a better recognition model while mitigating forgetting.
Figure A7: Performance of Models Using ICMscore without time penalty. (a) Top-1 accuracy comparison for image classification models trained on LM (blue), HM (green), and MM (yellow) images. The dashed line indicates the 20% chance level. (b) Continual learning performance showing Accuracy (left), BWT (middle), and Forgetting (right) for models using HM (green), MM (yellow), and LM (blue) replays. Error bars represent standard error over 5 independent runs. Evaluation metrics are detailed in Sec. 4.3.